# Don’t Waste It: Guiding Generative Recommenders with Structured Human Priors via Multi-head Decoding

### v2

@@ 1 Introduction
To navigate these multifaceted objectives, industrial recommendation systems have accumulated a wide array of post-hoc adjustments applied during the ranking or post-ranking stage Wang2020PersonalizedRF; McDonald2023Impatient; Wu2025GAS. We refer to this accumulated domain expertise as human priors. For example, diversity is often enforced by greedily selecting candidates that maximize a combined function of relevance and entropy (defined over manually tuned categories). To favor high-value interactions (e.g., purchases over clicks), practitioners typically build separate value models for each interaction type and apply heuristic weighting schemes. Similarly, balancing short-term engagement with long-term interests often involves temporal discounting heuristics or separate value models trained on different time horizons. Furthermore, ensuring adequate personalization for minority users frequently relies on first identifying these minority users and then optimizing a separate value model, or boosting content based on demographic features.
@@ 2 Related Work
We position our work at the intersection of generative recommendation, muti-interest and disentangled representation learning, and the integration of structured knowledge, motivated by the broader shift toward human-centered recommendation.
@@ 2.3 Integration of Human Priors and Structure
There is growing recognition that integrating structured, human-understandable knowledge, or human priors, can enhance model performance and interpretability.
@@ 3.2 Incorporating Human Priors via Conditioned Query Heads
This masking approach filters out all the incompatible items for the prior heads and ensures each head can focus exclusively on the subset of items aligned with its prior group. As a result, in contrast to the score in Eq. (1), the resulting score in Eq. (3) is tailored to different items with their prior information, which results in an explicit decomposition of user intent. Unlike conventional unsupervised approaches built on implicit latent factors Ma2019MacridVAE; guo2024DualVAE; Li2019MIND, our method allows for more model interpretability as it guarantees that the learned representation $\boldsymbol{q}_{k}$ is identifiable. In addition, while conventional approaches suffer from the inherent uncertainty arising from the entanglement of user preferences and their underlying latent factors, our method mitigates this issue by disentangling this complexity into a set of more tractable sub-tasks, thereby enhancing the computational efficiency.
@@ 3.6 Implementation Details
The proposed framework is model-agnostic and can be integrated with any generative recommender system that produces a dense user representation $\mathbf{h}_{T}$. The efficiency of the approach stems from the lightweight nature of the adapters [^2] We show each head only incurs an extra $0.14\%$ of overall parameters in Section 4.3. and the parallelizability of the query computations at the same depth level in the hierarchy, thus minimizing the latency overhead. Per-group indices $\{\Omega_{k}\}$ are pre-computed and cached, enabling efficient computation of the masked scores (Eq. (3)) via batched matrix multiplication.
@@ 4.1.2 MerRec (E-commerce)
MerRec LiMerRec2025 is derived from the Mercari C2C marketplace. It is characterized by exceptionally long interaction sequences (there are $119756$ users that have at least $2000$ interactions) and diverse user behaviors.
@@ 4.1.3 EB-NeRD (News)
Graph Prior: We construct an item co-engagement graph, where an edge exists if two items are interacted with by the same user. To discover underlying structural priors, for simplicity, we apply the off-the-shelf Leiden algorithm Traag2018FromLT from the igraph package, an established method for community detection that optimizes modularity and guarantees that the resulting communities are well connected. We control the influence of highly active users and merge very small clusters (details in Appendix 7.1). The resulting item clusters are used as graph-based priors, testing the framework’s ability to leverage community structures [^3] While this algorithmic clustering is a practical choice for our implementation, its success also underscores the framework’s tolerance for noisy or approximated priors, suggesting it can derive benefits even from imperfect structural information..
@@ 4.3 Main Results
Both ComiRec and REMI underperform the counterparts when we inject human prior into HSTU. The reason is that instead of relying on purely unsupervised methods to discover latent interests as in multi-interest networks, our approach uses these explicit priors as a form of weak supervision to guide the model in learning different user intents as well as disentangled representations. As for DualVAE, we observe it to be very unstable when implemented based on the deep HSTU backbone. With its original shallow encoders, it only marginally outperforms HSTU on Pixel8M and even slightly lags behind HSTU on MerRec and EB-NeRD.
@@ 4.4.2 User Interest Exploration
Exploration in recommender systems aims to uncover content a user may like but has not yet engaged with. Although it improves long-term user engagement, it is often believed to negatively affect near-term experience chenexplorationrs2021, similar to the exploration versus exploitation dilemma in reinforcement learning. We argue that adding multiple heads for human priors can somehow strike a good balance between the two. To evaluate this, we analyze the performance of our method on a targeted subset of users in the Pixel8M dataset using HLLM. Specifically, we identify users who interacted with item features during the forecast horizon of the test split, but who had never engaged with those features in their prior history. In these cases, the model’s success hinges on its ability to recommend items from entirely new categories. Applying this criterion, we identified a total of 283,497 such users.
@@
We observe that the relative improvements are consistently higher for this New Interest subset across all configurations. Moreover, incorporating Item Prior yields substantially more benefits compared to using only Temporal Prior (LT/ST). For example, HLLM with both LT/ST1 and Item Prior achieves a remarkable relative boost of +15.76% for NDCG@10, compared to only +3.58% for LT/ST1 alone. This supports the intuition that when learning the different categories together, the user embedding will be biased against the minority items, while dedicating specific heads to the minority items allows us to retain the capacity and encouraging exploration toward novel categories. Case studies in Section 8.1 further illustrate this insight.
@@ 4.5 Scalability
A key trend in recommender systems is the use of longer context lengths to enhance personalization. While scaling laws typically require more training data to improve larger models, our experimental setup with a fixed dataset introduces a trade-off: increasing the context length reduces the number of available training windows. To investigate this, we test HSTU with longer context lengths and larger model sizes on Pixel8M. In Figure 3, the left subplot shows that for size 4 and 1b, the base model (LT/ST1, equivalent to the discounted loss) struggles to benefit from increasing context as we increase the context length beyond 20 items. However, in the right subplot, when guided by human priors, the same model architecture continues to extract performance gains from longer contexts and larger model sizes, although the magnitude of increase slowly plateaus. This finding suggests that the structural information imposed by human priors facilitates more efficient learning, allowing the model to better benefit from increasing context and larger model sizes when the amount of training data is fixed.
@@ 4.7.2 Why it works: priors or just more heads?
Next, we show that the benefit of LT/ST interests is consistent across different model sizes of HSTU (Figure 6). Here, we train on $\tau=8$ using only the LT/ST prior. The baseline is only trained to predict the next one item, while LT/ST1 means that we only model one interest, and the objective reduces to the simple discounted loss over the next $\tau$ items. We see that explicitly assigning more different heads to model the long-term vs. short-term interests brings more improvements as we go from the baseline to $\text{seg}=2$, and then the benefit from further increasing the number of segments slowly diminishes. We observe a slight dip from size4 to 1b, and our hypothesis is that the number of target items used to train each head decreases when the number of interests increases. Nevertheless, the benefit of separately modeling the long-term and short-term interests is still significant at the scale of 1b, where the performance gain of the baseline model from increasing the model size has already flattens out at this scale.
@@ 5 Conclusion
By integrating human priors accumulated over post-hoc business rules directly into generative recommenders’ learning process, our proposed framework offers a principled approach to aligning recommendations with multifaceted human objectives. Our experiments demonstrate that prior-conditioned adapter heads not only enhance accuracy, but also improve diversity, novelty, and personalization, which are key dimensions of user experience that are often overlooked. The framework’s backbone-agnostic design and hierarchical composition strategy further enable flexible modeling of complex user intent, making it broadly applicable across scenarios.
@@ 6 Acknowledgment
We would like to thank Zhuokai Zhao, Benyu Zhang, Lizhu Zhang, Xiangjun Fan, and Abhishek Kumar for their support.
[…]
7 Experiment Details
7.1 Dataset Details and Preprocessing
Table 8: Statistics of the datasets after preprocessing. Filtering criteria vary slightly depending on the backbone model.
Dataset Backbone # Users # Items # Interactions
Pixel8M HSTU 561,737 398,261 57M
HLLM 2,220,506 404,182 102M
MerRec Both 119,754 1,255,665 521M
EB-NeRD Both 44,968 25,216 30M
Pixel8M. For HSTU, we filter out users with fewer than 50 interactions. For HLLM, we filter out users with fewer than 20 interactions.
  • Settings: For HSTU, we use a context length ($T$) of 50 and a forecast horizon ($\tau$) of 8. For HLLM, we use $T=10$ and $\tau=4$ due to its higher computation cost.
  • Modalities: Text inputs are “title”, “tag”, and “description”. Images are rescaled to $224\times 224$ pixels.
  • Item Prior Details: Over 5% of all items are labeled with the Miscellaneous tag, which appear to be unlabeled data that can randomly come from any of the other tags. We have to simply set all 8 binary features to be True for this tag. However, this also demonstrates the robustness of our method to noise in the human prior. The resulting normalized frequency distribution of the eight features is: Entertainment: 24.95%, Real life: 21.10%, Performance & Arts: 15.30%, Informational & Educational: 12.69%, Fictional character: 9.00%, Music: 6.29%, Gaming: 6.89%, and Science & technology: 3.77%. For example, the original tag “Food Production” is labeled with Real life, Informational & Educational, and Entertainment, while the original tag “Celebrities Mix” is labeled with Real life, Entertainment, Performance & Arts.
MerRec. We select the user subset with over 2,000 interactions.
  • Settings: For HSTU, we use a context length of $T=400$, and for HLLM, we use $T=50$, both to predict one item ahead.
  • Modalities: Text inputs are “c2_name” and “brand_name”.
EB-NeRD. We select the 44,968 users with over 512 interactions.
  • Settings: For HSTU, we use $T=50,\tau=8$. For HLLM, we use $T=24,\tau=4$.
  • Modalities: Text inputs are “title”, “subtitle”, and “topics”.
  • Graph Construction Details: For the co-engagement graph, we cap the contribution of highly active users by considering only their last 1000 interactions to prevent them from dominating the graph structure. When running the Leiden algorithm Traag2018FromLT (implemented via the igraph package), we use modularity as the optimization objective and tune the resolution parameter based on the desired number of clusters. Small clusters falling below a size threshold are merged into a single larger cluster.
7.2 Model Architectures and Hyperparameters
For both HSTU and HLLM backbones, we use a discount factor $\gamma=0.99$. However, we observed that a smaller $\gamma$ might lead to higher accuracy at the cost of lower diversity (Table 7).
HLLM Configurations. The choice of LLMs for HLLM depends on the available modalities in the dataset. For Pixel8M, which contains both text and images, we use Qwen2-VL-2B-Instruct as the Item LLM and Qwen2.5-1.5B as the User LLM. For MerRec and EB-NeRD, which only contain text, we use TinyLlama-1.1B-Chat-v1.0 as the Item LLM and TinyLlama_v1.1 as the User LLM.
HSTU Configurations. We experiment with various sizes [^5] For the 1B model, we use the same hyperparameters as TinyLlama as in the original HLLM paper. However, the total number of parameters is less than 1B due to the simplification of the feed-forward block in HSTU. of the HSTU model, detailed in Table 9. For all datasets, we remap product IDs based on the ones that still have been interacted with after filtering. We report the results based on size 3 unless noted.
A Specific Instantiation. We illustrate a concrete instantiation of our hierarchical framework (Section 3.3) using both Temporal and Graph Priors, corresponding to the configuration used on the EB-NeRD dataset (Figure 7). For visualization, we assume two graph clusters, $C_{1}$ (green) and $C_{2}$ (purple). In this example, the forecast horizon ($\tau=6$) is divided into two equal segments (LT/ST2), where $ST$ (short-term) is responsible for the first three targets and $LT$ (long-term) is responsible for the last three targets.
In the first layer (depth 1), the adapters specialize based on the temporal segments. The adapter $\mathcal{A}_{ST}^{(1)}$ is trained using only ground-truth items falling within the $ST$ segment, while $\mathcal{A}_{LT}^{(1)}$ is trained using items in the $LT$ segment. They output the intermediate representations $\mathbf{z}^{(1)}_{ST}$ and $\mathbf{z}^{(1)}_{LT}$, respectively.
Size 1 Size 2 Size 3 Size 4 1B
num_layers 4 8 12 16 22
num_heads 4 8 8 16 32
d_model 128 256 512 1024 2048
dropout 0.1 0.1 0.2 0.2 0.4
Total Params 12.4 M 26.8 M 64.3 M 181.6 M 658.9 M
Table 9: The list of hyperparameters in the five model sizes of HSTU, along with the total parameter counts.
Baseline Configurations. For ComiRec and REMI, we use the self-attention version, which demonstrates performance comparable to the dynamic routing version, but with stabler and faster training. We use a learning rate of $0.001$, search the number of interest from $\{1,2,4,8,16,32\}$, and report the best performance. For REMI, we tune the additional $\beta$ parameter for interest-aware hard negative mining from $\{0.1,1,4,10\}$ and set routing regularization weight $\lambda$ to $100$ following the original paper.
For DualVAE, we reimplement it under the window-wise sequential recommendation setting. We search the aspect number from $\{4,5,10\}$ and set the VAE latent dimension to $32$. We search dropout from $\{0.1,0.15,0.2\}$, $\gamma$ (for contrastive loss) and $\beta$ (for KL-divergence) from $\{1e-3,1e-2,1e-1,1\}$.
8 Additional Results
8.1 Case Studies
To illustrate the qualitative benefits of human priors, we present a case study in Figure 8 using the Pixel8M dataset. We analyze the predictions of the HLLM backbone, which leverages both text and image modalities [^6] Video descriptions are omitted in Figure 8 due to space constraints., configured with LT/ST2 and Item Prior, and we compare it against the baseline HLLM.
The user history (top two rows) does not contain videos related to “Informational & Educational”, but they indicate a latent user interest in military and geopolitics, since they include clips from the WWII movie Downfall, a game on evolution under nuclear waste, and a meme on “atomic egg explosion”.
The baseline HLLM (fourth row) struggles to synthesize this nuanced intent. Even after filtered to the “Informational & Educational” category, it produces relatively generic recommendations. In contrast, our proposed model demonstrates effective specialization (last row). The adapter head dedicated to the Long-Term interest within the “Informational & Educational” prior successfully captures the user’s latent interest, resulting in recommendations such as the revival of the Soviet Union and Russian territorial waters. The better recommendation quality is validated by the target items (third row). Starting from the second target item, the user engaged with content related to the Soviet Union, which was successfully predicted by our model, and the Second Sino-Japanese War. This shows that dedicating a specialized head to model underrepresented categories is effective at discovering the user’s hidden interests, which the baseline overlooked.
Figure 8: A case study on Pixel8M. Top two rows: browsing history. Third row: ground truth target items. Fourth row: HLLM baseline predictions (filtered to “Informational & Educational”). Last row: predictions from our model’s long-term “Informational & Educational” head.

### v3

@@ 1 Introduction
To navigate these multifaceted objectives, industrial recommendation systems have accumulated a wide array of post-hoc adjustments applied during the ranking or post-ranking stage McDonald2023Impatient; Wu2025GAS. We refer to this accumulated domain expertise as human priors. For example, diversity is often enforced by greedily selecting candidates that maximize a combined function of relevance and entropy (defined over manually tuned categories). To favor high-value interactions (e.g., purchases over clicks), practitioners typically build separate value models for each interaction type and apply heuristic weighting schemes pei2019value; de2023systematic. Similarly, balancing short-term engagement with long-term interests often involves temporal discounting heuristics or separate value models trained on different time horizons tang2019towards. Furthermore, ensuring adequate personalization for minority users frequently relies on first identifying these minority users and then optimizing a separate value model, or boosting content based on demographic features lin2025fair.
@@ 2 Related Work
We position our work at the intersection of generative recommendation, multi-interest and disentangled representation learning, and the integration of structured knowledge, motivated by the broader shift toward human-centered recommendation.
@@ 2.3 Integration of Human Priors and Structure
There is growing recognition that integrating structured, human-understandable knowledge, or human priors, can enhance model performance and interpretability (e.g., expert-defined in-domain taxonomy Yang2024Bespoke).
@@ 3.2 Incorporating Human Priors via Conditioned Query Heads
This masking approach filters out all the incompatible items for the prior heads and ensures each head can focus exclusively on the subset of items aligned with its prior group. As a result, in contrast to the score in Eq. (1), the resulting score in Eq. (3) is tailored to different items with their prior information, which leads to an explicit decomposition of user intent. Unlike conventional unsupervised approaches built on implicit latent factors Ma2019MacridVAE; guo2024DualVAE; Li2019MIND, our method allows for more model interpretability as it guarantees that the learned representation $\boldsymbol{q}_{k}$ is identifiable. In addition, while conventional approaches suffer from the inherent uncertainty arising from the entanglement of user preferences and their underlying latent factors, our method mitigates this issue by disentangling this complexity into a set of more tractable sub-tasks, thereby enhancing the computational efficiency.
@@ 3.6 Implementation Details
A Specific Instantiation. We illustrate a concrete instantiation of our hierarchical framework using both Temporal and Graph Priors, corresponding to the configuration used on the EB-NeRD dataset (Figure 1). For visualization, we assume two graph clusters, $C_{1}$ (green) and $C_{2}$ (purple). In this example, the forecast horizon ($\tau=6$) is divided into two equal segments (LT/ST2), where $ST$ (short-term) is responsible for the first three targets and $LT$ (long-term) is responsible for the last three targets.
In the first layer (depth 1), the adapters specialize based on the temporal segments. The adapter $\mathcal{A}_{ST}^{(1)}$ is trained using only ground-truth items within the $ST$ segment, while $\mathcal{A}_{LT}^{(1)}$ is trained using items in the $LT$ segment. They output the intermediate representations $\mathbf{z}^{(1)}_{ST}$ and $\mathbf{z}^{(1)}_{LT}$, respectively.
Efficiency. The proposed framework is model-agnostic and can be integrated with any generative recommender system that produces a dense user representation $\mathbf{h}_{T}$. The efficiency of the approach stems from the lightweight nature of the adapters [^2] We show each head only incurs an extra $0.14\%$ of overall parameters in Section 4.3. and the parallelizability of the query computations. Formally, let $K$ denote the total number of heads and $d$ the hidden dimension. The added parameter complexity is $O(K\cdot d^{2})$. The value of $K$ depends on the structure: it is the sum of categories (e.g., taxonomies), or the product for independent priors (e.g., the number of temporal segments multiplied with the number of graph clusters). During inference, adapters at the same depth level are executed in parallel, ensuring the latency overhead remains low. Per-group indices $\{\Omega_{k}\}$ are pre-computed and cached, enabling efficient computation of the masked scores (Eq. (3)) via batched matrix multiplication.
@@ 4.1.2 MerRec (E-commerce)
MerRec LiMerRec2025 is derived from the Mercari C2C marketplace. It is characterized by exceptionally long interaction sequences (there are $119756$ users who have at least $2000$ interactions) and diverse user behaviors.
@@ 4.1.3 EB-NeRD (News)
Graph Prior: We construct an item co-engagement graph, where an edge exists if two items are interacted with by the same user. To discover underlying structural priors, for simplicity, we apply the off-the-shelf Leiden algorithm Traag2018FromLT from the igraph package, an established method for community detection that optimizes modularity and guarantees that the resulting communities are well connected. We control the influence of highly active users and merge very small clusters (details in Appendix 7.1). The resulting item clusters are used as graph-based priors, testing the framework’s ability to leverage community structures [^3] While this algorithmic clustering is a practical choice for our implementation, its success also underscores the framework’s tolerance to noisy or approximated priors, suggesting it can derive benefits even from imperfect structural information..
@@ 4.3 Main Results
Both ComiRec and REMI underperform their counterparts when we inject human prior into HSTU. The reason is that instead of relying on purely unsupervised methods to discover latent interests as in multi-interest networks, our approach uses these explicit priors as a form of weak supervision to guide the model in learning different user intents as well as disentangled representations. As for DualVAE, we observe it to be very unstable when implemented based on the deep HSTU backbone. With its original shallow encoders, it only marginally outperforms HSTU on Pixel8M and even slightly lags behind HSTU on MerRec and EB-NeRD.
@@ 4.4.2 User Interest Exploration
Exploration in recommender systems aims to uncover content a user may like but has not yet engaged with. Although it improves long-term user engagement, it is often believed to negatively affect near-term experience chenexplorationrs2021, similar to the exploration versus exploitation dilemma in reinforcement learning. We argue that adding multiple heads for human priors can strike a good balance between the two. To evaluate this, we analyze the performance of our method on a targeted subset of users in the Pixel8M dataset using HLLM. Specifically, we identify users who interacted with item features during the forecast horizon of the test split, but who had never engaged with those features in their prior history. In these cases, the model’s success hinges on its ability to recommend items from entirely new categories. Applying this criterion, we identified a total of 283,497 such users.
@@
We observe that the relative improvements are consistently higher for this New Interest subset across all configurations. Moreover, incorporating Item Prior yields substantially more benefits compared to using only Temporal Prior (LT/ST). For example, HLLM with both LT/ST1 and Item Prior achieves a remarkable relative boost of +15.76% for NDCG@10, compared to only +3.58% for LT/ST1 alone. This supports the intuition that when learning the different categories together, the user embedding will be biased against the minority items, while dedicating specific heads to the minority items allows us to retain the capacity and encourage exploration toward novel categories. Case studies in Section 8 further illustrate this insight.
@@ 4.5 Scalability
A key trend in recommender systems is the use of longer context lengths to enhance personalization zhang2026efficient. While scaling laws typically require more training data to improve larger models, our experimental setup with a fixed dataset introduces a trade-off: increasing the context length reduces the number of available training windows. To investigate this, we test HSTU with longer context lengths and larger model sizes on Pixel8M. In Figure 4, the left subplot shows that for size 4 and 1b, the base model (LT/ST1, equivalent to the discounted loss) struggles to benefit from increasing context as we increase the context length beyond 20 items. However, in the right subplot, when guided by human priors, the same model architecture continues to extract performance gains from longer contexts and larger model sizes, although the magnitude of increase slowly plateaus. This finding suggests that the structural information imposed by human priors facilitates more efficient learning, allowing the model to better benefit from increasing context and larger model sizes when the amount of training data is fixed.
@@ 4.7.2 Why it works: priors or just more heads?
Next, we show that the benefit of LT/ST interests is consistent across different model sizes of HSTU (Figure 7). Here, we train on $\tau=8$ using only the LT/ST prior. The baseline is only trained to predict the next one item, while LT/ST1 means that we only model one interest, and the objective reduces to the simple discounted loss over the next $\tau$ items. We see that explicitly assigning more different heads to model the long-term vs. short-term interests brings more improvements as we go from the baseline to $\text{seg}=2$, and then the benefit from further increasing the number of segments slowly diminishes. We observe a slight dip from size4 to 1b, and our hypothesis is that the number of target items used to train each head decreases when the number of interests increases. Nevertheless, the benefit of separately modeling the long-term and short-term interests is still significant at the scale of 1b, where the performance gain of the baseline model from increasing the model size has already flattened out at this scale.
@@ 5 Conclusion
By integrating human priors accumulated over post-hoc business rules directly into generative recommenders’ learning process, our proposed framework offers a principled approach to aligning recommendations with multifaceted human objectives. We view the capacity to leverage human priors not as a dependency, but as a strategic advantage tailored to the unique landscape of recommender systems. Unlike domains where human priors are scarce, the recommendation field possesses decades of market-validated expertise, resulting in many high-quality priors. Our experiments demonstrate that prior-conditioned adapter heads not only enhance accuracy, but also improve diversity, novelty, and personalization, which are key dimensions of the user experience that are often overlooked. The framework’s backbone-agnostic design and hierarchical composition strategy further enable flexible modeling of complex user intent, making it broadly applicable across scenarios.
@@ 6 Acknowledgment
We would like to thank Zhuokai Zhao and Lizhu Zhang for their support.